OpenStudy (anonymous):

A grain silo is shown below. What is the volume of grain that could completely fill this silo rounded to the nearest whole number? Use for pi.

OpenStudy (anonymous):

how do i put the picture on here?

OpenStudy (saifoo.khan):

Use the "attach file" option.

OpenStudy (btaylor):

Even though your pic didn't post (try replying again), \[V =\pi r^{2}h\]

OpenStudy (btaylor):

OK then, to add half the sphere: \[V = 2/3 pi r^{3}\]

OpenStudy (anonymous):

ok how bout the rest of it

OpenStudy (saifoo.khan):

volume of cylinder + volume of semi-circle

OpenStudy (anonymous):

can you substitute my numbers in the equation

OpenStudy (saifoo.khan):

semi sphere*

OpenStudy (saifoo.khan):

do you know the formula for each? @helpppleaseee7

OpenStudy (anonymous):

no i dont

OpenStudy (btaylor):

We just gave them to you. \[V = \pi r^{2}h+ 2/3 \pi r^{3}\]

OpenStudy (anonymous):

can you substitute the numbers from my problem into there?

OpenStudy (anonymous):

@BTaylor and @saifoo.khan

OpenStudy (btaylor):

\[V=168\times6^{2} \pi +2/3\times6^{3}\pi\] \[V=6048\pi + 144\pi = 6192\pi\] \[6192\pi = 19453 ft^{3}\]

OpenStudy (anonymous):

thanks man i understand now!